A bird flies for 4 sec with a velocity of |t-2| m/s in a straight line where
t=time in seconds .It covers a distance of

Grade:12

5 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Pratik

given

V= |t-2| 

dx/dt = |t-2|

dx = |t-2|dt

oxdx  = o4|t-2|dt

 x  = o2(2-t)dt  + 24(t-2)dt

      =4

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Pratik Sharma
14 years ago

how the step 5 come ?

Jatin
6 years ago
givenV= |t-2| dx/dt = |t-2|dx = |t-2|dto∫xdx = o∫4|t-2|dt x = o∫2(2-t)dt + 2∫4(t-2)dt =4givenV= |t-2| dx/dt = |t-2|dx = |t-2|dto∫xdx = o∫4|t-2|dt x = o∫2(2-t)dt + 2∫4(t-2)dt =4givenV= |t-2| dx/dt = |t-2|dx = |t-2|dto∫xdx = o∫4|t-2|dt x = o∫2(2-t)dt + 2∫4(t-2)dt =4
Mohit poonia
6 years ago
As we know that 
At t=0  u=2 m/s     and at t=2  v=0 m/s
Therefore by v=u+at   we get,
0=2+a2   =)     a=-1m/s2
And by s=vt-1/2at2  we get 
S= 0×2-1/2×-1×2×2     =)     s=2m ---------------1
So the first part is complete and now the second part starts
We know that at  t=2seconds   u=0m/s     and at     t=4sec   v=2m/s so total time =2sec
So by v=u+at
2=0+a×2       =)         a=1m/s2  
Now by  s=ut+1/2at2      
s= 0×2+1/2×1×2×2      =)         s=2m      -------------2
So by adding 1 and 2
We get distance travelled by bird =2+2=4 meters
And thats the answer for those who find the calculus in this question difficult.  And also feel free to ask any doubts involved in this solution.
 
 
ankit singh
3 years ago
 
At t=0  u=2 m/s     and at t=2  v=0 m/s
Therefore by v=u+at   we get,
0=2+a2   =)     a=-1m/s2
And by s=vt-1/2at2  we get 
S= 0×2-1/2×-1×2×2     =)     s=2m ---------------1
So the first part is complete and now the second part starts
We know that at  t=2seconds   u=0m/s     and at     t=4sec   v=2m/s so total time =2sec
So by v=u+at
2=0+a×2       =)         a=1m/s2  
Now by  s=ut+1/2at2      
s= 0×2+1/2×1×2×2      =)         s=2m      -------------2
So by adding 1 and 2
We get distance travelled by bird =2+2=4 meters